# Superfiltering: Weak-to-Strong Data Filtering for Fast Instruction-Tuning

### v1

@@ Abstract
Instruction tuning is critical to improve LLMs but usually suffers from low-quality and redundant data. Data filtering for instruction tuning has proved important in improving both the efficiency and performance of the tuning process. But it also leads to extra cost and computation due to the involvement of LLMs in this process. To reduce the filtering cost, we study Superfiltering: Can we use a smaller and weaker model to select data for finetuning a larger and stronger model? Despite the performance gap between weak and strong language models, we find their highly consistent capability to perceive instruction difficulty and data selection results. This enables us to use a much smaller and more efficient model to filter the instruction data used to train a larger language model. Not only does it largely speed up the data filtering, but the filtered-data-finetuned LLM achieves even better performance on standard benchmarks. Extensive experiments validate the efficacy and efficiency of our approach. Our codes, models, and data will be released at https://github.com/tianyi-lab/Superfiltering.
@@
Ming Li1, Yong Zhang2, Shwai He1, Zhitao Li2, Hongyu Zhao1, Jianzong Wang2, Ning Cheng2, Tianyi Zhou1thanks: Corresponding author 1University of Maryland 2Ping An Technology (Shenzhen) Co., Ltd. minglii@umd.edu, jzwang@188.com, tianyi@umd.edu
@@ 1 Introduction
The landscape of natural language processing has witnessed a transformative change with the introduction of large language models (LLMs) like GPT-3 Brown et al. (2020), GPT-4OpenAI (2023), LLaMA Touvron et al. (2023a, b), Mistral Jiang et al. (2023), etc Penedo et al. (2023); Scao et al. (2022). These models offer unprecedented capabilities in generating contextually rich, coherent, and often creative text. A key advantage of them is the ability to follow instructions, which leads to promising performance on zero-shot (prompting) or few-shot (in-context learning) tasks. To achieve it, a supervised learning technique named instruction tuning Wei et al. (2022); Longpre et al. (2023a) is commonly utilized, which finetunes an LLM to produce preferred responses on a broad range of tasks described by natural language instructions.
Earlier works of instruction tuning focus on creating large, varied, and high-quality datasets of various tasks with responses curated by human experts Khashabi et al. (2020); Ye et al. (2021); Wei et al. (2022); Wang et al. (2022); Du et al. (2022), which can be bottlenecked by the intensive human labor. An alternative is to generate the data by a powerful teacher LLM Wang et al. (2023b); Taori et al. (2023); Xu et al. (2023); Li et al. (2023a) but the quality is hard to control and largely depends on the teacher. LIMA Zhou et al. (2023) finds that a mere 1,000 human-crafted high-quality data could significantly improve an LLM’s instruction-following capability, based on which they posit that LLMs acquire most knowledge during the pretraining and thus a few data suffices for instruction tuning.
To further free the human labor in data curation and accelerate the instruction tuning process, a line of recent works apply an extra filter algorithm to select data from the existing dataset. However, the model used in the filtering process usually needs to be as powerful as ChatGPT Chen et al. (2023b); Lu et al. (2023) or requires additional reward data training Du et al. (2023); Bukharin and Zhao (2023), or is the student model itself Li et al. (2023b), which leads to additional expensive cost and latency due to the inference on these large filter models, especially when the original dataset is large while only a tiny fraction of data needs to be selected. These paradigms are presented in Figure 1 (a) and (b). To reduce the filtering cost, we study Superfiltering: Can we use a smaller and weaker model as a filter to select instruction-tuning data for training a larger and stronger model? This was first studied for training small classification models by Coleman et al. (2020) while the effectiveness on the open-domain instruction dataset is un-explored. Recently, Weak-to-Strong Generalization Burns et al. (2023) proposes to utilize a weaker ChatGPT to generate data used to finetune a stronger GPT4 model, which shares a similar spirit with our Superfiltering as depicted in Figure 1(c).
@@
  • Weak-to-Strong Consistency on Data Filtering: We discover a strong consistency between small and large LLMs in perceiving and evaluating the difficulty of instruction tuning data.
@@ 2.2 Formulation and Motivations
Dataset Model(weak-to-strong) Rank Correlation $\uparrow$ Overlap Ratios $\uparrow$
Name Size Perplexity IFD score 5% 10% 15%
Alpaca GPT-2 124M 0.726 0.679 0.28 0.41 0.49
GPT-2-large 774M 0.790 0.682 0.26 0.40 0.50
GPT-2-XL 1.5B 0.802 0.693 0.27 0.40 0.49
GPT-NEO 1.3B 0.846 0.802 0.38 0.51 0.59
LLaMA2-7B 7B 1.000 1.000 1.00 1.00 1.00
Alpaca-GPT4 GPT-2 124M 0.730 0.788 0.24 0.40 0.51
GPT-2-large 774M 0.795 0.820 0.21 0.36 0.48
GPT-2-XL 1.5B 0.800 0.818 0.18 0.33 0.45
GPT-NEO 1.3B 0.842 0.876 0.33 0.52 0.62
LLaMA2-7B 7B 1.000 1.000 1.00 1.00 1.00
WizardLM 70k GPT-2 124M 0.763 0.802 0.42 0.54 0.61
GPT-2-large 774M 0.809 0.848 0.44 0.58 0.65
GPT-2-XL 1.5B 0.821 0.855 0.44 0.57 0.65
GPT-NEO 1.3B 0.857 0.893 0.52 0.63 0.69
LLaMA2-7B 7B 1.000 1.000 1.00 1.00 1.00
@@ 4.2 Implementation Details
4.3 Evaluation Metrics and Benchmarks
4.3.1 Pair-wise comparison
Evaluating responses generated by Large Language Models (LLMs) like GPT-4 remains a complex and ongoing research area, particularly for open-domain questions where establishing a clear ground truth is challenging. Traditional methods often fall short in assessing the instruction-following ability of these models. Recent trends, however, involve using LLMs themselves, such as GPT-4, as evaluators, a practice that has gained widespread acceptance in the field Touvron et al. (2023b); Chiang et al. (2023); Dettmers et al. (2023); Liu et al. (2023). Previous studies Zheng et al. (2023); Li et al. (2023c); Sottana et al. (2023) have shown that GPT4’s evaluations are consistent with human evaluations. We utilized the testing instruction set from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023) which contain $218$ and $80$ diverse human-curated instructions respectively.
4.3.2 AlapcaEval Leaderboard
4.3.3 Open LLM Leaderboard
@@ 5 Experimental Result
Dataset/ Superfilter Pairwise normal-$\uparrow$ Huggingface Open LLM Leaderboard $\uparrow$ AlpacaEval normal-$\uparrow$
@@ 5.1 Main results
In this section, we present the evaluation results of three different evaluation settings as described in the previous section as shown in Table 2. The Pair-Wise Winning Score indicates the result directly comparing with the corresponding model trained with full data. These values that are greater than $1.0$ represent better responses generated by our Superfiltering models than full data models. The detailed win-tie-lose numbers are presented in Figure 2. Moreover, the performance of our models and baseline models on the Huggingface Open LLM Leaderboard and the AlpacaEval Leaderboard are also presented in Table 2 where we can see our models using $5\%$, $10\%$, $15\%$ data outperform the models trained with full data on both benchmarks on both LLaMA2-7B and LLaMA-13B settings. These results further showcase the effectiveness of our automatically selected data. Moreover, the usefulness of Superfiltering on the high-quality Alpaca-GPT4 dataset further shows the potential of our method, which is astonishing that a pure statistical metric based on a weak language model like GPT-2 is able to filter the responses generated by GPT-4.
@@ 6 Further Discussion
In the realm of data selection for language model instruction tuning, our Superfiltering introduces a transformative advantage: the unnecessity of training for even weak language models. Traditional proxy-based methods like Coleman et al. (2020) and Nguyen et al. (2022) are required to further train weak models to bridge the performance gap with stronger models. However, our study reveals that pre-trained weak models are naturally effectively capable of acting as the proxies for strong models when utilizing the IFD for data selection, without requiring any additional training.
Moreover, in the context of instruction tuning data selection, model training is always necessary if no extra strong models like ChatGPT or other trained reward models are utilized. Lu et al. (2023) utilizes chatGPT to tag the instruction datasets and train LLMs for tagging instruction samples based on these data. Wei et al. (2023) first splits a range of subsets from the original data and records each fine-tuned model’s performance on the validation set as the labels of data quality. Then a self-attention network as the data selector is trained. Cao et al. (2023) utilizes a trained regression model to estimate the inference losses as data qualities on several datasets.
The performances of the resulting efficient selection models are appealing while a possible concern is the generalizability since they all need extra performance indicators, i.e. development set, in existing datasets. On the contrary, our approach directly utilizes established, widely accepted models like GPT-2, which are known for their broad applicability and generalization ability. These models do not necessitate fine-tuning on specific datasets, thereby reducing the risk of out-of-distribution issues. Our method demonstrates that these smaller models can be deployed in a “plug-and-play” manner, achieving commendable performance immediately. This innovative approach not only simplifies the data selection process but also revolutionizes the efficiency and applicability of such methods in large language model instruction tuning.
@@ 7 Related Work
7.1 Instruction Tuning
Recent advancements in natural language processing (NLP) have been significantly influenced by instruction tuning, a method that tailors large language models (LLMs) for varied tasks using explicit instructions Wei et al. (2022); Sanh et al. (2022); Longpre et al. (2023a). This approach has enhanced LLMs’ ability to understand and follow instructions in diverse contexts. Initial research in this area focused on expanding dataset sizes to improve instruction-following capabilities Honovich et al. (2023); Wang et al. (2023b). Additionally, innovative methods, such as using LLMs to generate instructional data, are being explored to streamline and enhance the instruction tuning process Wang et al. (2023b); Xu et al. (2023); He et al. (2023); Li et al. (2023a).
@@ 7.2 Instruction Tuning Data Selection
To further select the data for more efficient instruction tuning, existing automatic data selection methods mainly utilize extra LLMs for the selection. Lu et al. (2023) utilizes proprietary chatGPT to tag the instruction data to ensure diversity and complexity. Chen et al. (2023b) utilizes proprietary LLMs chatGPT and Claude2 to assess the quality of the instruction data, generating both ratings and explanations. Du et al. (2023) and Bukharin and Zhao (2023) utilize an extra reward model to assess the quality of data and utilize these scores as a part of their method. Li et al. (2023b) firstly proposes a self-guided method in which no extra LLMs are utilized but still needs to calculate Instruction-Following Difficulty (IFD) scores based on the original pre-trained LLM. Though effective, these methods overly rely on large language models and are too time-consuming to put into practical use.
@@ 7.3 Small Model Proxies for Large Models
The use of proxy models is increasingly recognized in machine learning, particularly when resources are constrained or there is a limited understanding of the original model’s architecture. Chen et al. (2023a) and Hase et al. (2020) demonstrate the utility of lightweight proxy models in evaluating free-text rationales. Similarly, Puigcerver et al. (2021) leverages embeddings from expert models with a k-nearest neighbors classifier to simplify the training of more complex systems. Coleman et al. (2020) and the FAMIE Nguyen et al. (2022) apply downscaled proxy models in fields like image classification and information extraction, utilizing techniques such as layer removal and knowledge distillation for aligning these proxies with larger models. Building on this, Burns et al. (2023) explores the concept of enhancing larger models through weak supervision, training them on labels generated by weaker models. By extending the "Weak to Strong" concept to LLM instruction tuning data selection, our research employs pre-trained smaller models as proxies, demonstrating their effectiveness in assessing instruction complexity, thereby bridging the gap between the comprehensive capabilities of larger models and the agility of smaller ones.
@@ 8 Conclusion
This paper presented “Superfiltering”, a novel and efficient approach for data filtering in the instruction tuning of LLMs. By effectively utilizing weaker models as proxies for evaluating instructional data, particularly in the context of IFD scores, we achieved a significant leap in efficiency, accelerating the data filtering process largely. The experimental results affirm that our method considerably reduces computational overhead while maintaining or even improving the instructional capabilities of LLMs. Thus, Superfiltering stands as a testament to our initial hypothesis and objectives, marking a substantial contribution to the field of natural language processing by offering a scalable, resource-efficient, and effective strategy for the advancement of AI technologies.

### v2

@@ Abstract
Instruction tuning is critical to improve LLMs but usually suffers from low-quality and redundant data. Data filtering for instruction tuning has proved important in improving both the efficiency and performance of the tuning process. But it also leads to extra cost and computation due to the involvement of LLMs in this process. To reduce the filtering cost, we study Superfiltering: Can we use a smaller and weaker model to select data for finetuning a larger and stronger model? Despite the performance gap between weak and strong language models, we find their highly consistent capability to perceive instruction difficulty and data selection results. This enables us to use a much smaller and more efficient model to filter the instruction data used to train a larger language model. Not only does it largely speed up the data filtering, but the filtered-data-finetuned LLM achieves even better performance on standard benchmarks. Extensive experiments validate the efficacy and efficiency of our approach.
@@
Ming Li1, Yong Zhang2, Shwai He1, Zhitao Li2, Hongyu Zhao1, Jianzong Wang2, Ning Cheng2, Tianyi Zhou1thanks: Corresponding author 1University of Maryland 2Ping An Technology (Shenzhen) Co., Ltd. minglii@umd.edu, jzwang@188.com, tianyi@umd.edu Project: https://github.com/tianyi-lab/Superfiltering
@@ 1 Introduction
Earlier works of instruction tuning on Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a); Jiang et al. (2023) focus on creating large, varied, and high-quality datasets of various tasks with responses curated by human experts Khashabi et al. (2020); Ye et al. (2021); Wei et al. (2022); Wang et al. (2022); Du et al. (2022), which can be bottlenecked by the intensive human labor. An alternative is to generate the data by a powerful teacher LLM Wang et al. (2023b); Taori et al. (2023); Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); Xu et al. (2024) but the quality is hard to control and largely depends on the teacher. LIMA Zhou et al. (2023) finds that a mere 1,000 human-crafted high-quality data could significantly improve an LLM’s instruction-following capability, based on which they posit that LLMs acquire most knowledge during the pretraining and thus a few data suffices for instruction tuning.
To further free the human labor in data curation and accelerate the instruction tuning process, a line of recent works apply an extra filter algorithm to select data from the existing dataset. However, the model used in the filtering process usually needs to be as powerful as ChatGPT Chen et al. (2023b); Lu et al. (2023) or requires additional reward data training Du et al. (2023); Bukharin and Zhao (2023), or is the student model itself Li et al. (2023b, 2024b), which leads to additional expensive cost and latency due to the inference on these large filter models, especially when the original dataset is large while only a tiny fraction of data needs to be selected. These paradigms are presented in Figure 1 (a) and (b). To reduce the filtering cost, we study Superfiltering: Can we use a smaller and weaker model as a filter to select instruction-tuning data for training a larger and stronger model? This was first studied for training small classification models by Coleman et al. (2020) while the effectiveness on the open-domain instruction dataset is un-explored. Recently, Weak-to-Strong Generalization Burns et al. (2023) proposes to utilize a weaker ChatGPT to generate data used to finetune a stronger GPT4 model, which shares a similar spirit with our Superfiltering as depicted in Figure 1(c).
@@
  • Weak-to-Strong Consistency on Data Filtering: We reveal the strong consistency between small and large LLMs in perceiving and evaluating the difficulty of instruction tuning data.
@@ 2.2 Formulation and Motivations
Dataset Model Rank Correlation $\uparrow$ Overlap Ratios $\uparrow$
Name Perplexity IFD score 5% 10% 15%
Alpaca GPT-2 0.726 0.679 0.28 0.41 0.49
GPT-2-large 0.790 0.682 0.26 0.40 0.50
GPT-2-XL 0.802 0.693 0.27 0.40 0.49
GPT-NEO 0.846 0.802 0.38 0.51 0.59
LLaMA2-7B 1.000 1.000 1.00 1.00 1.00
Alpaca-GPT4 GPT-2 0.730 0.788 0.24 0.40 0.51
GPT-2-large 0.795 0.820 0.21 0.36 0.48
GPT-2-XL 0.800 0.818 0.18 0.33 0.45
GPT-NEO 0.842 0.876 0.33 0.52 0.62
LLaMA2-7B 1.000 1.000 1.00 1.00 1.00
Wizard 70k GPT-2 0.763 0.802 0.42 0.54 0.61
GPT-2-large 0.809 0.848 0.44 0.58 0.65
GPT-2-XL 0.821 0.855 0.44 0.57 0.65
GPT-NEO 0.857 0.893 0.52 0.63 0.69
LLaMA2-7B 1.000 1.000 1.00 1.00 1.00
@@ 4.2 Implementation Details
4.3 Evaluation Metrics
4.3.1 Automatic Evaluation
To evaluate the effectiveness of our method, we utilize 3 commonly used automatic evaluation metrics, including (1) Pair-wise Comparison, (2) Open LLM Leaderboard and (3) Alpaca Eval.  [^1] Detailed description of evaluation metrics can be found in Appendix B.
4.3.2 Human Evaluation
To further validate the effectiveness of our method, we conducted a further human study to evaluate the effectiveness of our method. Specifically, we randomly sample 100 instructions from the WizardLM test set to form a new instruction set. Then $3$ human participants are given the task of comparing the responses generated by the comparing models with the criteria same as the previous pair-wise evaluation, i.e. Helpfulness, Relevance, Accuracy, and Level of Detail. For each comparison, 3 options are given (Win, Tie, and Loss) and the final results are determined by the majority voting of the participants. The human studies are conducted based on LLaMA2-7B on two models: (1) Alpaca 5% VS. Alpaca 100%. (2) Alpaca-GPT4 5% VS. Alpaca-GPT4 100%.
@@ 5 Experimental Result
Dataset/ Superfilter Pairwise $\uparrow$ Huggingface Open LLM Leaderboard $\uparrow$ AlpacaEval $\uparrow$
@@ 5.1 Main results
In this section, we present the evaluation results of three different evaluation settings as described in the previous section as shown in Table 5. The Pair-Wise Winning Score indicates the result directly comparing with the corresponding model trained with full data. These values that are greater than $1.0$ represent better responses generated by our Superfiltering models than full data models. The detailed win-tie-lose numbers are presented in Figure 2. Moreover, the performance of our models and baseline models on the Huggingface Open LLM Leaderboard and the AlpacaEval Leaderboard are also presented in Table 5 where we can see our models using $5\%$, $10\%$, $15\%$ data outperform the models trained with full data on both benchmarks on both LLaMA2-7B and LLaMA-13B settings. These results further showcase the effectiveness of our Superfiltering. Moreover, the usefulness of Superfiltering on the high-quality Alpaca-GPT4 dataset further shows the potential of our method, which is astonishing that a weak language model like GPT-2 is able to filter the data responses generated by GPT-4.
For the human evaluation, we compare the performances between models trained with 5% Superfiltring data and full data based on LLaMA2-7B models, on Alpaca and Alpaca-GPT4 Datasets. In the comparison (1) Alpaca 5% VS. Alpaca 100%, our model wins on $50$ out of $100$ instruction, ties on $18$, and losses on $32$ instructions. In the comparison (2) Alpaca-GPT4 5% VS. Alpaca-GPT4 100%, our model wins on $49$ out of $100$ instruction, ties on $5$, and losses on $46$ instructions. This human evaluation of Superfiltering further validates our method.
@@
5.2 Comparison with Other Methods
In this subsection, we compare our method with three other widely accepted instruction-tuning data selection methods on LLaMA2-7B using the Alpaca Dataset, in terms of performance and efficiency. “ChatGPT score” represents utilizing ChatGPT to evaluate the quality of the Alpaca data samples proposed by Chen et al. (2023b). “Reward-model score” represents utilizing extra reward models to rate the given data samples, in this experiment, “OpenAssistant/reward-model-deberta-v3-large-v2” is utilized following Bukharin and Zhao (2023); Du et al. (2023). “IFD score” represents directly utilizing the base model to be trained to calculate the Instruction-Following Difficulty scores proposed by Li et al. (2023b).
Comparison Pairwise Winning Score $\uparrow$ Time
[…]
Superfiltering (ours) - - - 8
vs. ChatGPT score 1.028 1.174 1.170 120
vs. Reward score 1.280 1.096 1.147 1400
vs. IFD score (LLaMA2-7B) 0.853 0.761 0.927 161
Table 4: Comparison with Other Methods in terms of performance and efficiency. The Pairwise Winning Scores are calculated between models using our method and other methods. All the comparisons are performed by GPT-4 on the WizardLM test set and the values that are greater than $1.0$ represent our models are better and vice versa. The time shown represents the time used for data filtering.
As shown in the Table 4. Superfiltering outperforms ChatGPT and Reward-model scores in terms of winning scores and filtering efficiency. Compared with Superfiltering, utilizing ChatGPT for selection needs further costs on utilizing API models, and its efficiency is largely constrained by the rate limits set by the API company. Utilizing outside reward models is a cost-free method, while the efficiency can not be guaranteed: it is 175×175$\times$175 × slower than our Superfiltering. Directly calculating IFD scores based on the base model is a promising method as it is the only method that has a better performance than our Superfiltering, as it selects the data that best matches the model. However, our Superfilrtering is 20×20$\times$20 × faster. From the comparison, it is observed that our model largely reduces the time used for data filtering, and we are the only method that makes the filtering time shorter than the later training time, making this process feasible in practical usage. Moreover, our filtering method can be implemented on consumer-level GPUs with much smaller graphic memories like 6 GB since we only utilize the GPT-2 model, while all other methods require large industry-level GPUs.
@@ 6 Further Discussion
Our Superfiltering introduces a transformative advantage: the unnecessity of training for even weak language models and the unnecessity of the extra hold-out sets.
Traditional proxy-based methods like Coleman et al. (2020) and Nguyen et al. (2022) are required to further train weak models to bridge the performance gap with stronger models. In the context of instruction tuning data selection, model training or a hold-out set is always necessary if no extra strong models like ChatGPT or other trained reward models are utilized. Lu et al. (2023) utilizes chatGPT to tag the instruction datasets and train LLMs for tagging instruction samples based on these data. Li et al. (2023d) requires a holdout test set as the indicator for the data improvement. Cao et al. (2023) proposes to train several models based on losses on unseen datasets and obtain a regression model to estimate data qualities.
However, our study reveals that pre-trained weak models are naturally effectively capable of acting as proxies for strong models when utilizing the IFD for data selection. These models do not necessitate fine-tuning on specific datasets, thereby reducing the risk of out-of-distribution issues. This innovative approach not only simplifies the data selection process but also revolutionizes the efficiency and applicability of such methods in large language model instruction tuning.
6.2 Sperfiltering as Dataset Assessment
Our finding, the consistency of perplexity-based metrics across weak and strong language models, might push forward the usage of small language models as proxy models, not only in the area of data selection. Moreover, our finding provides an efficient and general way to assess the instruction data of the whole dataset.
As illustrated in Figure 3, a preliminary assessment of the dataset can be made by examining the violin plots of the Instruction Followed Difficulty (IFD) scores. The classic Alpaca dataset, known for its relatively lower quality, exhibits a violin plot with wide upper and lower sections, indicating significant variance in data quality. The WizardLM dataset, generated by prompting ChatGPT 3.5, displays complex instructions but is affected by generation noise, reflected by its violin plot which is wide at the top and has a long, narrow tail. In contrast, the Alpaca-GPT4 dataset, created using GPT-4, demonstrates a higher quality lower bound, as evidenced by its violin plot with a more compressed lower section, indicating fewer low-quality instructions. To our knowledge, there is no visualization that can represent the quality of a dataset with different characteristics, (some work might draw the distribution of verb-noun pairs, but it only represents the diversity and can not reflect the overall quality of the dataset).
6.3 Why Weak-to-Strong Consistent?
We hypothesize this Weak-to-Strong consistency can be explained from (1) the perspective of Perplexity and (2) the perspective of String Pattern, presented in Appendix D.
6.4 Superfiltering with Diversity
We further incorporate the diversity metric into the Superfiltering, which further reduces the data used for training, presented in the Appenix A.
@@ 7.1 Instruction Tuning Data Selection
Instruction tuning Wei et al. (2022); Sanh et al. (2022); Longpre et al. (2023a); Liu et al. (2023a) is a widely-used training paradigm to equip LLMs with the instruction-following ability. To further select the data for more efficient instruction tuning, existing automatic data selection methods mainly utilize extra LLMs for the selection. Lu et al. (2023) utilizes proprietary chatGPT to tag the instruction data to ensure diversity and complexity. Chen et al. (2023b) utilizes proprietary LLMs chatGPT and Claude2 to assess the quality of the instruction data, generating both ratings and explanations. Du et al. (2023) and Bukharin and Zhao (2023) utilize an extra reward model to assess the quality of data and utilize these scores as a part of their method. Li et al. (2023b) firstly proposes a self-guided method in which no extra LLMs are utilized but still needs to calculate Instruction-Following Difficulty (IFD) scores based on the original pre-trained LLM. Though effective, these methods overly rely on large language models and are too time-consuming to put into practical use.
@@ 7.2 Small Model Proxies for Large Models
The use of proxy models is increasingly recognized in machine learning, particularly when resources are constrained or there is a limited understanding of the original model’s architecture. Chen et al. (2023a) and Hase et al. (2020) demonstrate the utility of lightweight proxy models in evaluating free-text rationales. Similarly, Puigcerver et al. (2021) leverages embeddings from expert models with a k-nearest neighbors classifier to simplify the training of more complex systems. Coleman et al. (2020) and the FAMIE Nguyen et al. (2022) apply downscaled proxy models in fields like image classification and information extraction, utilizing techniques such as layer removal and knowledge distillation for aligning these proxies with larger models. Burns et al. (2023) explores the concept of enhancing larger models through weak supervision, and training on weaker model labels.
@@ 8 Conclusion
In this paper, we reveal the consistency between weak and strong language models in perceiving instruction difficulties. Based on the consistency, we present “Superfiltering”, a novel and efficient approach for data filtering in the instruction tuning of LLMs. By effectively utilizing weaker models as proxies for evaluating instructional data, we achieve a significant leap in efficiency, accelerating the data filtering process largely. The experimental results affirm that our method considerably reduces computational overhead while maintaining or even improving the instruction tuning performance of LLMs. It is astonishing that a weak language model like GPT-2 is able to filter the data responses generated by GPT-4. Our Superfiltering marks a substantial contribution by offering a scalable, resource-efficient, and effective strategy for the advancement of AI technologies.
Limitations
Our study introduces “Superfiltering” as an innovative approach to enhance the instruction tuning process of LLMs through efficient data selection using smaller models. While we have observed promising results, several areas warrant further exploration: (1) Superfiltering primarily utilizes IFD scores to select data, focusing on instruction difficulty. Future enhancements should include additional dimensions such as data diversity. (2) There are $3$ datasets and $2$ LLaMA2 base models are involved, further research is recommended to explore its effectiveness with a broader array of LLMs and Datasets.
Appendix A Superfiltering with Diversity
Though Superfiltering shows promising performance with great efficiency, it can also be implemented together with other dimensions such as data diversity. Motivated by recent work that further includes Diversity metrics in the data selection process, we introduce an extended version of Superfiltering, Superfiltering with Diversity (Superfiltering.D).
We hypothesize that the diversity metrics work better when implemented on a high-quality data subset than the whole dataset with mixed quality. Thus we propose to first utilize Superfiltering to select a subset with relatively high quality, then further utilize Facility Location Function  [^2] https://apricot-select.readthedocs.io/en/latest/functions/facilityLocation.html to further compress the selected data. Compared with other diversity metrics, the Facility Location Function can strike a balance between capturing diversity and ensuring the representation of different clusters or regions within the data, it ensures a global view of the given high-quality subset.
To further preserve the efficiency of our Superfiltering.D, we utilize “sentence-transformers/all-MiniLM-L6-v2” Reimers and Gurevych (2019) as the encoder, which only has approximately 80M parameters. In our preliminary experiments on the Alpaca and Alpaca-GPT4 dataset, where we first select $20\%$ of the data by Superfiltering, then utilize the Facility Location Function to further select $2\%$ of the data. The models trained with $2\%$ of the data have a comparable or better performance than full data models.
Dataset/ Superfiltering Pairwise $\uparrow$ Huggingface Open LLM Leaderboard $\uparrow$
Base Model Ratio(Size) Winning Score Average ARC HellaSwag MMLU TruthfulQA
Alpaca/ 100% 1.000 55.25 54.35 78.65 47.02 40.98
LLaMA2-7B 2%(1,040) 1.028 55.43 55.97 79.89 45.51 40.34
Alpaca-GPT4/ 100% 1.000 58.71 54.69 80.05 47.89 52.21
LLaMA2-7B 2%(1,040) 1.078 58.70 56.48 80.30 45.23 52.79
Table 5: Comparison between Superfiltering.D with 2% data and full data model on two datasets (Alpaca and Alpaca-GPT4). The finetuned models are evaluated by the pair-wise winning score (comparison to the baseline model finetuned on 100% data) and Open LLM Leaderboard. In the parathesis are the ratio of data being used and its exact number.
Appendix B Evaluation Metric
B.1 Pair-wise comparison
Evaluating responses generated by Large Language Models (LLMs) like GPT-4 remains a complex and ongoing research area, particularly for open-domain questions where establishing a clear ground truth is challenging. Traditional methods often fall short in assessing the instruction-following ability of these models. Recent trends, however, involve using LLMs themselves, such as GPT-4, as evaluators, a practice that has gained widespread acceptance in the field Touvron et al. (2023b); Chiang et al. (2023); Dettmers et al. (2023); Liu et al. (2023b). Previous studies Zheng et al. (2023); Li et al. (2023c); Sottana et al. (2023) have shown that GPT4’s evaluations are consistent with human evaluations. We utilized the testing instruction set from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023) which contain $218$ and $80$ diverse human-curated instructions respectively.
B.2 AlapcaEval Leaderboard
B.3 Open LLM Leaderboard
Appendix C Prompt for Evaluation
The detailed pair-wise comparison prompt for the pair-wise comparison is in Figure 4.
  Prompt for Performance Evaluation
  System Prompt
You are a helpful and precise assistant for checking the quality of the answer.
User Prompt
[Question]
Question
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
Figure 4: The prompt we used to request ChatGPT or GPT4 to evaluate the responses.
Appendix D Why Weak-to-Strong Consistent?
In this section, we provide the hypothesis on why this Weak-to-Strong consistency exists on two perspectives.
From the perspective of Perplexity: The consistency between IFD scores is built on the consistency of perplexities since it is the ratio between two kinds of perplexities. As for the reason why the rankings of perplexities are consistent between small and large language models, we hypothesize that both small and large language models are trained on similar corpus (almost all the existing data corpus), which means they’re exposed to the same underlying distribution of language. Even though larger models can capture more nuances and details due to their increased parameter count, both sizes of models are ultimately trying to approximate the same underlying language structure.
From the perspective of String Pattern: We also examined verb-noun pairs of the data samples filtered by IFD scores on both weak (GPT-2) and strong (LLaMA2-7B) language models on the Alapca dataset and found a strong consistency in these verb-noun pairs. The results are listed in Table 6. From the table, we can find out that the language models tend to assign higher IFD scores to those data that require a lot of creativity, thinking skills, and deep understanding while assigning lower IFD scores to those data that are more about following rules and need less creativity. This holds true for both small and large language models.
LLaMA2-7B (Top 5% IFD) GPT2 (Top 5% IFD)
Verb Noun Count Verb Noun Count
write story 119 write story 117
generate list 104 generate story 84
generate story 89 create story 52
write essay 61 generate list 47
create list 44 write essay 39
write post 41 create list 29
write article 39 write paragraph 26
create story 36 write post 25
generate recipe 34 generate recipe 19
create recipe 31 give example 17
LLaMA2-7B (Least 5% IFD) GPT2 (Least 5% IFD)
Verb Noun Count Verb Noun Count
rewrite sentence 140 rewrite sentence 124
edit sentence 56 edit sentence 74
classify sentence 51 classify sentence 30
change sentence 42 replace word 24
convert sentence 33 change sentence 23
arrange word 29 construct query 20
rearrange word 27 convert sentence 20
categorize sentence 26 arrange word 17
find word 24 find word 17
classify text 18 classify animal 16
Table 6: Verb-noun pairs of the top and least 5% IFD scores calculated by LLaMA2-7B and GPT2 on the Alpaca dataset.